# Enhancing Graph U-Nets for Mesh-Agnostic Spatio-Temporal Flow Prediction

### v1

@@ Abstract
This study aims to overcome the conventional deep-learning approaches based on convolutional neural networks, whose applicability to complex geometries and unstructured meshes is limited due to their inherent mesh dependency. We propose novel approaches to improve mesh-agnostic spatio-temporal prediction of transient flow fields using graph U-Nets, enabling accurate prediction on diverse mesh configurations. Key enhancements to the graph U-Net architecture, including the Gaussian mixture model convolutional operator and noise injection approaches, provide increased flexibility in modeling node dynamics: the former reduces prediction error by 95% compared to conventional convolutional operators, while the latter improves long-term prediction robustness, resulting in an error reduction of 86%. We also investigate transductive and inductive-learning perspectives of graph U-Nets with proposed improvements. In the transductive setting, they effectively predict quantities for unseen nodes within the trained graph. In the inductive setting, they successfully perform in mesh scenarios with different vortex-shedding periods, showing 98% improvement in predicting the future flow fields compared to a model trained without the inductive settings. It is found that graph U-Nets without pooling operations, i.e. without reducing and restoring the node dimensionality of the graph data, perform better in inductive settings due to their ability to learn from the detailed structure of each graph. Meanwhile, we also discover that the choice of normalization technique significantly impacts graph U-Net performance. In conclusion, the proposed enhancements to the graph U-Net architecture and the insights gained from transductive and inductive-learning perspectives lay the foundation for developing versatile mesh-agnostic models that enable accurate and reliable spatio-temporal flow predictions across a wide range of real-world applications.
@@
journal: Journal of Computational Physics
@@ 1 Introduction
Instantaneous transient flow prediction is essential for the realization of digital twins, which serve as the cornerstone for real-time forecasting and decision making in a variety of engineering fields. More specifically, the integration of accurate unsteady flow prediction models is critical to the fidelity and effectiveness of digital twins, enabling them to mirror their physical counterparts in real-time. However, conventional high-fidelity computational fluid dynamics (CFD) analysis, which is one of the most useful approaches for understanding fluid-flow behavior, has the disadvantage of requiring significant computational resources for practical applications [1, 2].
@@
In this context, artificial intelligence (AI) is emerging as a transformative solution that plays a critical role in predicting real-time flow fields within digital-twin scenarios [3, 4, 5, 6, 7, 8]. Harnessing the power of AI can significantly reduce the computational burden while improving the accuracy of transient flow simulations and automating the entire analysis process. In addition, AI-driven approaches offer the potential for real-time prediction, which is essential for digital twins that can adapt to dynamic flow conditions. These advances pave the way for more sophisticated and responsive digital twins, ensuring that their digital counterparts can operate in real-time with unprecedented accuracy and efficiency, revolutionizing predictive maintenance and operational effectiveness across CFD-related industries.
@@
Particularly, convolutional neural networks (CNNs) have led to significant advances in the prediction of both steady and unsteady flow fields [9, 10, 11, 12, 13]. While CNNs have introduced more efficient and physically meaningful flow-prediction algorithms due to their spatial locality property, they are inherently designed for structured grids and therefore face significant limitations when confronted with unstructured meshes. This limitation is exacerbated as unstructured meshes become more prevalent in real-world industrial sites due to their flexibility in handling complex geometries and dynamic mesh scenarios. Consequently, the reliance on CNNs, which are not applicable to unstructured meshes, highlights a significant limitation of their potential within digital twins for real-world industrial scenarios. This gap highlights the need for advanced deep-learning models that can handle the various unstructured grid systems for generality and versatility in practical applications.
@@
In that regard, graph neural networks (GNNs) can be a powerful solution, effectively bridging the gap left by CNNs through their inherent mesh scalability [6, 14, 15, 16, 17, 15, 18, 19]. They can handle different structures of mesh datasets by learning directly from the graph consisting of nodes and edges without the spatial topology restrictions (structured grid) inherent in CNNs [20, 21]. Therefore, GNNs are perfectly positioned to address the challenges associated with both structured and unstructured meshes commonly encountered in CFD simulations. The mesh flexibility of GNNs enables them to seamlessly adapt to various mesh configurations during training and inference, making them highly compatible with diverse CFD datasets. By directly operating on graph-structured data, GNNs can effectively capture the intricate spatial relationships and dependencies present in the mesh, allowing them to model the complex dynamics and evolution of unsteady flow fields with high fidelity.
@@
Among the various approaches based on GNNs, graph U-Nets, which originated from the U-Net architecture, have been introduced for their performance in node/graph classification tasks [22]. U-Nets have been widely adopted for their efficiency in feature extraction and localization across different resolution scales, enabling precise segmentation and reconstruction tasks [23, 24, 25, 26, 27]. This ability of the U-Nets to capture multi-scale features is particularly well-suited for flow-prediction tasks where flow structures with a wide range of length scales exist. By integrating the U-Net framework with GNNs, graph U-Nets inherit these strengths, enabling exceptional performance even in untrained mesh scenarios [22]. However, their comprehensive validation has not been conducted in regression tasks with high-dimensional spatio-temporal output space, which are the most common in practical flow problems. Therefore, this work attempts to explore the potential of this combination, U-Nets and GNNs, which would ensure the multi-scale nature of U-Nets even in the unstructured-data situation, allowing them to capture both local and global flow patterns and make them highly effective in predicting transient flow fields with complex spatio-temporal dependencies.
@@
This study focuses on leveraging the graph U-Net architecture to predict vortex-shedding flow over a circular cylinder, which is one of the most famous benchmark flow problems for spatio-temporal prediction [28]. We propose several novel approaches to augment the synergy between the strengths of GNNs in capturing mesh-based dependencies and the spatio-temporal dynamics of transient flows. The proposed methods introduce several key enhancements to the traditional graph U-Net architecture, including the Gaussian mixture model (GMM) convolutional operator [29] for increased flexibility in modeling node dynamics, and noise injection approaches [30] to mitigate error accumulation over long time-stepping. Throughout this study, extensive parametric studies are conducted to investigate the impact of various architectural choices and hyperparameters on the performance of the proposed graph U-Net model for unsteady flow-field prediction.
@@
    The first study to explore the potential of graph U-Nets in unsteady flow-field prediction and to fill the gap in the existing literature, which primarily focuses on CNN-based U-Net approaches that rely on the use of structured CFD grids.
@@
    A comprehensive parametric study investigating the impact of various architectural choices and hyperparameters of graph U-Nets on their performance in unsteady flow-field prediction is conducted, serving as a valuable guide for future researchers.
@@
    Gaussian mixture model (GMM) convolutional operator is proposed to be applied instead of the traditional graph convolutional network (GCN) due to its greater flexibility in modeling the complex flow dynamics between neighboring nodes.
@@
    The incorporation of noise-injection approaches to mitigate error accumulation over long rollouts (or time-stepping) is proposed, ensuring the dramatic improvement in temporal prediction accuracy.
@@
    Both transductive and inductive performances of graph U-Nets are investigated, demonstrating their predictive capabilities on both untrained nodes and untrained meshes.
@@
    To enable inductive learning in graph U-Nets, we extend the training scenario by leveraging meshes with different vortex shedding periods. Furthermore, we demonstrate that omitting the pooling operation is crucial for accurate predictions in this setting.
  7. 7.
    Meanwhile, we also discover that the choice of normalization technique significantly affects the inductive performance of graph U-Net. The superiority of the best normalization approach was found to be due to the temporal consistency of the meshes, emphasizing the need to select normalization methods based on the characteristics of the meshes used.
By addressing these key aspects and introducing novel enhancements to the graph U-Net architecture, our study presents a comprehensive and tailored approach for mesh-agnostic spatio-temporal forecasting of unsteady flow fields. We believe that the proposed framework not only advances the state of the art in graph-based deep learning for fluid dynamics, but also lays the foundation for the development of more accurate, efficient, and versatile digital twins in a variety of engineering domains.
@@ 2.1.1 Graph convolutional network (GCN) operator
It is important to note that GCNs are inspired by the foundational principles of CNNs, specifically their ability to perform feature extraction through the convolution operation [31]. While CNNs apply convolutional filters over regular, grid-like data structures (e.g., images), GCNs adapt this concept to non-Euclidean graph-structured data where the layout is irregular. Both methods aggregate local neighborhood information to capture spatial hierarchies, but GCNs extend this idea to nodes and their connections, enabling feature extraction from graphs in a manner analogous to how CNNs operate on images. The conventional GCN operation in node-wise formulation is as follows:
@@ 4.3.2 Noise injection approaches
Our findings offer insights into the application of noise-injection techniques in graph-based models for fluid-dynamics simulations. Both the I-noise and I/O-noise approaches significantly enhance the model’s robustness to error accumulation during rollouts. This study highlights the importance of selecting an appropriate noise-injection approach and carefully tuning the noise size based on the specific problem and data scale to optimize the performance of graph U-Nets in unsteady flow prediction tasks.
@@ 5.1 Transductive-learning settings
From the various aspects investigated in this section, it can be concluded that the graph U-Nets with the proposed settings in this study (with GMM operator, with the selected best pooling ratio, and with the noise-injection approach) can perform successfully in transductive learning for unsteady flow-field prediction. The trained models effectively predict quantities for unseen nodes within the same graph used during training by leveraging the trained information obtained from the truncated graph structure. This highlights the generalizability of graph U-Nets in capturing the dynamics of transient flow fields, even when dealing with unseen nodes in familiar meshes.
@@ 6 Conclusions and future work
This study presented a comprehensive investigation of graph U-Nets specialized for unsteady flow-field prediction, focusing on their mesh-agnostic spatio-temporal forecasting capabilities. By introducing key enhancements to the traditional graph U-Net architecture, such as the GMM convolutional operator and noise injection approaches, we have demonstrated the potential of graph U-Nets to accurately predict complex fluid dynamics across diverse mesh configurations and flow conditions.
@@
    GMM convolutional operator is proposed to be applied instead of the traditional GCN due to its greater flexibility in modeling the flow dynamics between neighboring nodes: GMM shows 95% improvement over GCN operators in terms of MSE.
@@
    Noise-injection approaches to mitigate error accumulation over long rollouts are incorporated. We found that these techniques lead to around 86% reduction in temporal-prediction error, provided that the appropriate noise size is chosen.
@@
    We investigated both the transductive and inductive performance of graph U-Nets, demonstrating their predictive capabilities on both trained and untrained meshes. For transductive learning, we discovered that the graph U-Net can predict the quantities on the unseen nodes with satisfactory accuracy even when trained on a mesh with the rear half of the domain truncated. For inductive learning, we observed that 98% improvement can be obtained in predicting future flow fields for a specific mesh scenario by applying the inductive-learning settings compared to a model trained without the inductive settings.
@@
The proposed graph U-Net framework has the potential to revolutionize the field of digital twins through its versatile and flexible application in unsteady flow-field prediction. By accurately and efficiently predicting unsteady flow fields across various mesh configurations and generalizing to different geometries and flow conditions as verified in this study, graph U-Nets can significantly improve the fidelity, responsiveness, and adaptability of digital twins. Although this study demonstrates the effectiveness of graph U-Nets using a benchmark vortex shedding problem, the methodology’s potential goes far beyond this single application. As the field of graph-based deep learning continues to evolve, we expect graph U-Nets to play an increasingly crucial role in enabling accurate and efficient predictions of complex fluid-dynamics phenomena.
Future work may include extending the graph U-Net architecture to be robust for much longer rollout predictions by applying various temporal numerical schemes or deep-learning architectures specialized for temporal predictions [43, 44]. For example, applying the forward Euler method with graph U-Nets that predict $\delta u$ (which is the amount of change in velocity $u$) instead of predicting $u$ directly can be an option. Additionally, the training algorithm can be modified with multi-step rollout [45], which uses the loss function computed from the next several snapshots all at once, instead of using only the immediately next snapshot. Finally, graph-based prediction will be coupled with the uncertainty-quantification approach [46] and then extended to the Bayesian design optimization applications, where optimal aerodynamic designs can be obtained through an adaptive sampling (also known as active learning) procedure that takes into account the predictive uncertainty of the trained AI models [47, 48].

### v2

@@ Abstract
This study aims to overcome the limitations of conventional deep-learning approaches based on convolutional neural networks in complex geometries and unstructured meshes by exploring the potential of Graph U-Nets for unsteady flow-field prediction. We present a comprehensive investigation of Graph U-Nets, originally developed for classification tasks, now tailored for mesh-agnostic spatio-temporal forecasting of fluid dynamics. Our focus is on enhancing their performance through systematic hyperparameter tuning and architectural modifications. We propose novel approaches to improve mesh-agnostic spatio-temporal prediction of transient flow fields using Graph U-Nets, enabling accurate prediction on diverse mesh configurations. Key enhancements to the Graph U-Net architecture, including the Gaussian-mixture-model convolutional operator and noise injection approaches, provide increased flexibility in modeling node dynamics: the former reduces prediction error by 95% compared to conventional convolutional operators, while the latter improves long-term prediction robustness, resulting in an error reduction of 86%. We demonstrate the effectiveness of these enhancements in both transductive and inductive learning settings, showcasing the adaptability of Graph U-Nets to various flow conditions and mesh structures. This work contributes to the field of reduced-order modeling for computational fluid dynamics by establishing Graph U-Nets as a viable and flexible alternative to convolutional neural networks, capable of accurately and efficiently predicting complex fluid flow phenomena across diverse scenarios.
@@
journal: Expert Systems with Applications
@@ 1 Introduction
Instantaneous transient flow prediction is essential for the realization of digital twins, which serve as the cornerstone for real-time forecasting and decision making in a variety of engineering fields. More specifically, the integration of accurate unsteady flow prediction models is critical to the fidelity and effectiveness of digital twins, enabling them to mirror their physical counterparts in real-time. However, conventional high-fidelity computational fluid dynamics (CFD) analysis, which is one of the most useful approaches for understanding fluid-flow behavior, has the limitation of requiring significant computational resources for practical applications [1, 2].
@@
In this context, artificial intelligence (AI) approaches have emerged as promising alternatives, playing a critical role in predicting real-time flow fields within digital-twin scenarios [3, 4, 5, 6, 7, 8]. Leveraging the power of AI can significantly reduce the computational burden while improving the accuracy of transient flow simulations and automating the entire analysis process. In addition, AI-driven approaches offer the potential for real-time prediction, which is essential for digital twins that can adapt to dynamic flow conditions. These advances pave the way for more sophisticated and responsive digital twins, ensuring that their digital counterparts can operate in real-time with unprecedented accuracy and efficiency, revolutionizing predictive maintenance and operational effectiveness across CFD-related industries.
@@
Particularly, convolutional neural networks (CNNs) have led to significant advances in the prediction of both steady and unsteady flow fields [9, 10, 11, 12, 13, 14]. While CNNs have introduced more efficient and physically meaningful flow-prediction algorithms due to their spatial locality property, they are inherently designed for structured grids and therefore face significant limitations when confronted with unstructured meshes. This limitation is exacerbated as unstructured meshes become more prevalent in real-world industrial sites due to their flexibility in handling complex geometries and dynamic mesh scenarios. Consequently, the reliance on CNNs, which are not applicable to unstructured meshes, highlights a significant limitation of their potential within digital twins for real-world industrial scenarios. This gap highlights the need for advanced deep-learning models that can handle the various unstructured grid systems for generality and versatility in practical applications.
@@
In that regard, graph neural networks (GNNs) can be a powerful solution, effectively bridging the gap left by CNNs through their inherent mesh scalability [6, 15, 16, 17, 18, 16, 19, 20, 21]. They can handle different structures of mesh datasets by learning directly from the graph consisting of nodes and edges without the spatial topology restrictions (structured grid) inherent in CNNs [22]. Therefore, GNNs are perfectly positioned to address the challenges associated with both structured and unstructured meshes commonly encountered in CFD simulations. The mesh flexibility of GNNs enables them to seamlessly adapt to various mesh configurations during training and inference, making them highly compatible with diverse CFD datasets. By directly operating on graph-structured data, GNNs can effectively capture the intricate spatial relationships and dependencies present in the mesh, allowing them to model the complex dynamics and evolution of unsteady flow fields with high fidelity.
@@
Among the various approaches based on GNNs, Graph U-Nets, which originated from the U-Net architecture, have been introduced for their performance in node/graph classification tasks [23]. U-Nets have been widely adopted for their efficiency in feature extraction and localization across different resolution scales, enabling precise segmentation and reconstruction tasks [24, 25, 26, 27, 28, 29, 30, 31]. This ability of the U-Nets to capture multi-scale features is particularly well-suited for flow-prediction tasks where flow structures with a wide range of length scales exist. By integrating the U-Net framework with GNNs, Graph U-Nets inherit these strengths, enabling exceptional performance even in untrained mesh scenarios [23]. However, their comprehensive validation has not been conducted in regression tasks with high-dimensional spatio-temporal output space, which are the most common in practical flow problems. Therefore, this work attempts to explore the potential of this combination, U-Nets and GNNs, which would ensure the multi-scale nature of U-Nets even in the unstructured-data situation, allowing them to capture both local and global flow patterns and make them highly effective in predicting transient flow fields with complex spatio-temporal dependencies.
@@
This study focuses on leveraging the Graph U-Net architecture to predict vortex-shedding flow over a circular cylinder, which is one of the most famous benchmark flow problems for spatio-temporal prediction [32]. We propose several novel approaches to augment the synergy between the strengths of GNNs in capturing mesh-based dependencies and the spatio-temporal dynamics of transient flows. The proposed methods introduce several key enhancements to the traditional Graph U-Net architecture, including the Gaussian mixture model (GMM) convolutional operator [33] for increased flexibility in modeling node dynamics, and noise injection approaches [34] to mitigate error accumulation over long time-stepping. Throughout this study, extensive parametric studies are conducted to investigate the impact of various architectural choices and hyperparameters on the performance of the Graph U-Net model in order to optimize its performance specifically for fluid dynamics applications.
@@
    A comprehensive exploration of Graph U-Nets in unsteady flow-field prediction, bridging the gap between structured grid-based CNN approaches and the need for flexible, mesh-agnostic real-time prediction models in CFD domain.
@@
    Systematic investigation and optimization of Graph U-Net architecture for fluid dynamics applications, including:
    1. a)
      Implementation of the Gaussian mixture model (GMM) convolutional operator, demonstrating increased flexibility and accuracy in modeling node dynamics.
    2. b)
      Analysis of pooling strategies, revealing their impact on prediction accuracy and generalization capabilities.
    3. c)
      Implementation of noise injection techniques to significantly enhance long-term prediction stability.
    4. d)
      Comparative assessment of normalization methods, uncovering their crucial role in model performance across different mesh scenarios.
@@
    In-depth investigation of both transductive and inductive learning performance, demonstrating the ability of Graph U-Nets to generalize to untrained node regions and entirely new mesh configurations.
@@
    Novel insights into inductive learning for Graph U-Nets, including:
    1. a)
      Successful extension of the training scenario to encompass meshes with varying vortex shedding periods.
    2. b)
      The counter-intuitive finding that omitting pooling operations leads to superior performance in inductive settings.
    3. c)
      Discovery of the significant impact of normalization techniques on inductive learning, with performance highly dependent on whether the mesh structure remains static or changes dynamically over time.
@@
    Demonstration of substantial improvements in prediction accuracy and long-term stability through carefully tuned Graph U-Net models, establishing their viability for practical fluid dynamics applications.
@@
    Development of a comprehensive framework for adapting and optimizing Graph U-Nets, originally designed primarily for classification tasks, to perform effectively in regression tasks for spatio-temporal flow prediction. This work demonstrates the versatility of Graph U-Nets beyond their conventional classification applications, unlocking new potential for their use in complex flow dynamics predictions.
By addressing these key aspects and introducing novel enhancements to the Graph U-Net architecture, our study presents a comprehensive and tailored approach for mesh-agnostic spatio-temporal forecasting of unsteady flow fields. We believe that this work not only advances the application of graph-based models in computational fluid dynamics but also provides valuable insights for researchers and practitioners seeking to leverage Graph U-Nets as an efficient, accurate, and adaptable method for flow field prediction across a variety of engineering domains.
@@ 2.1.1 Graph convolutional network (GCN) operator
It is important to note that GCNs are inspired by the foundational principles of CNNs, specifically their ability to perform feature extraction through the convolution operation [35, 36]. While CNNs apply convolutional filters over regular, grid-like data structures (e.g., images), GCNs adapt this concept to non-Euclidean graph-structured data where the layout is irregular. Both methods aggregate local neighborhood information to capture spatial hierarchies, but GCNs extend this idea to nodes and their connections, enabling feature extraction from graphs in a manner analogous to how CNNs operate on images. The conventional GCN operation in node-wise formulation is as follows:
@@ 4.3.2 Noise injection approaches
Our findings offer insights into the application of noise injection techniques in graph-based models for fluid-dynamics simulations. Both the I-noise and I/O-noise approaches significantly enhance the model’s robustness to error accumulation during rollouts. This study highlights the importance of selecting an appropriate noise injection approach and carefully tuning the noise size based on the specific problem and data scale to optimize the performance of Graph U-Nets in unsteady flow prediction tasks.
@@ 5.1 Transductive-learning settings
From the various aspects investigated in this section, it can be concluded that the Graph U-Nets with the proposed settings in this study (with GMM operator, with the selected best pooling ratio, and with the noise injection approach) can perform successfully in transductive learning for unsteady flow-field prediction. The trained models effectively predict quantities for unseen nodes within the same Graph Used during training by leveraging the trained information obtained from the truncated graph structure. This highlights the generalizability of Graph U-Nets in capturing the dynamics of transient flow fields, even when dealing with unseen nodes in familiar meshes.
@@ 6 Conclusions and future work
This study presented a comprehensive investigation of Graph U-Nets specialized for unsteady flow-field prediction, focusing on their mesh-agnostic spatio-temporal forecasting capabilities. By introducing key enhancements to the traditional Graph U-Net architecture, such as the GMM convolutional operator and noise injection approaches, we have established a framework for adapting these models, originally developed primarily for classification tasks, to perform effectively in regression tasks for spatio-temporal flow prediction across diverse mesh configurations and flow conditions.
@@
    The GMM convolutional operator is proposed to be applied instead of the traditional GCN due to its greater flexibility in modeling the flow dynamics between neighboring nodes: GMM shows 95% improvement over GCN operators in terms of MSE.
@@
    Noise injection approaches to mitigate error accumulation over long rollouts are incorporated. We found that these techniques lead to around 86% reduction in temporal-prediction error, provided that the appropriate noise size is chosen.
@@
    We investigated both the transductive and inductive performance of Graph U-Nets, demonstrating their predictive capabilities on both trained and untrained meshes. For transductive learning, we discovered that the Graph U-Net can predict the flow quantities on the unseen nodes with satisfactory accuracy even when trained on a mesh with the rear half of the domain truncated. For inductive learning, we observed that 98% improvement can be obtained in predicting future flow fields for a specific mesh scenario by applying the inductive-learning settings compared to a model trained without the inductive settings.
@@
The proposed Graph U-Net framework has the potential to advance the field of digital twins through its application in real-time unsteady flow-field prediction based on unstructured grids, highlighting the versatility of Graph U-Nets beyond their conventional classification applications. By accurately and efficiently predicting unsteady flow fields across various mesh configurations and generalizing to different geometries and flow conditions as verified in this study, Graph U-Nets can significantly improve the fidelity, responsiveness, and adaptability of digital twins. Although this study demonstrates the effectiveness of Graph U-Nets using a benchmark vortex-shedding problem, the methodology’s potential goes far beyond this single application. As the field of graph-based deep learning continues to evolve, we expect Graph U-Nets to play an increasingly crucial role in enabling accurate and efficient predictions of complex fluid-dynamics phenomena.
Future work will include extending the Graph U-Net architecture to be robust for much longer rollout predictions by applying various temporal numerical schemes or deep-learning architectures specialized for temporal predictions [48, 49]. For example, applying the forward Euler method with Graph U-Nets that predict $\delta u$ (which is the amount of change in velocity $u$) instead of predicting $u$ directly can be an option. Additionally, the training algorithm can be modified with multi-step rollout [50], which uses the loss function computed from the next several snapshots all at once, instead of using only the immediately next snapshot. Finally, graph-based prediction will be coupled with the uncertainty-quantification approach [51, 52] and then extended to the Bayesian-design-optimization applications, where optimal aerodynamic designs can be obtained through an adaptive-sampling (also known as active learning) procedure that takes into account the predictive uncertainty of the trained AI models [53, 54].